# QAPyramid: Fine-grained Evaluation of Content Selection for Text Summarization

### v1

@@ Appendix A Complementary Results
A.1 Length and Repetition Penalty
@@
Table 5: The pearson correlation between $p^{l}_{i}*\text{QAPyramid{}}$ and effective summary length with different $\alpha$s.
A.2 System Scores
BART PEGASUS BRIO BRIO-Ext MatchSum Llama-3-8B-Inst. Llama-3-70B-Ins. Mixtral-8$\times$7B-Inst. Mixtral-8$\times$22B-Inst. GPT-4
@@ A.2 System Scores
Appendix B Examples, Prompts, and Annotation UIs
@@
Reference: Loeb says he filed the lawsuit and doesn’t want want money from his "ex" Nick Loeb reportedly wants to prevent Vergara from destroying the embryos . Vergara spoke of freezing embryos with Loeb in a 2013 interview .
Summary (generated by MatchSum): The 42-year-old actress and star of the hit TV sitcom "Modern Family" split from businessman Nick Loeb in May 2014. Loeb is suing the Colombian-born actress in Los Angeles to prevent Vergara from destroying their two embryos conceived through in vitro fertilization in November 2013, according to published reports by New York Daily News and In Touch magazine.
@@
What did someone say? he filed the lawsuit and does n’t want want money from his " ex "
@@
Who doesn’t someone want something from? his " ex "
@@
QA to Statement Convert the question and answer into a statement. Start your answer with "Statement:"

### v2

@@ Appendix A Complementary Results
A.1 Correlations with Other Annotation Benchmarks
We evaluate AutoQAPyramid on the SummEval benchmark (Fabbri et al., 2021a). Since SummEval provides multiple reference summaries, we test both single-reference and multi-reference variants of our metric, correlating their outputs against the benchmark’s human relevance scores. The multi-reference variants computes the score by taking the average across triplets extracted from all references. The results, presented in Table 4, show that the multi-reference variant outperforms the single-reference one in five out of six cases.
Despite this, the overall correlations with human judgments are lower than expected. We hypothesize this is due to a protocol mismatch: our metric is reference-based, but SummEval’s annotations are reference-free (annotators were not shown a reference summary). Such a discrepancy is known to depress correlations, with Liu et al. (2023b) finding a system-level Pearson’s correlation of -0.25 between reference-based and reference-free human judgments. Indeed, when we compare our reference-based metric with another reference-based metric, ROUGE-1, we observe a strong Pearson correlation of 0.73. This confirms our metric is performing correctly but is misaligned with SummEval’s reference-free setup.
This finding highlights a critical gap in existing resources and validates the need for our newly collected human annotations, which provide a benchmark specifically designed to assess fine-grained, reference-aware relevance between summaries.
System-level Summary-level
r $\rho$ $\tau$ r $\rho$ $\tau$
AutoQAPyramid single-ref 0.27 0.02 -0.03 0.06 0.05 0.04
AutoQAPyramid multi-ref 0.26 0.12 0.07 0.10 0.10 0.07
Table 4: Pearson (r), Spearman ($\rho$), and Kendall ($\tau$) correlations of AutoQAPyramid on SummEval.
A.2 Length and Repetition Penalty
@@
[…]
@@
A.3 System Scores
We show the system scores in Table 6.
Appendix B Annotation Details
B.1 QA Generation Annotation
We collected human-written QA pairs on Amazon Mechanical Turk. QA pairs were collected for 500 reference summaries from CNN/DM. For each summary, we broke it into sentences. And for each sentence, we found all the predicates (verbs). There are 7.6 verbs per summary on average. So it’s 3800 verbs in total. For each job (HIT), we presented the annotator with one verb highlighted in one sentence (Figure 6), and the annotator only needed to write QA pairs for this verb (on average 2.2 pairs per verb). It typically takes 1-2 minutes to finish one HIT. All 31 annotators we hired had gone through qualification tests and training. They knew the task pretty well. So, in total, it took about 3800 * 2 mins = 7600 minutes = 127 hours (about 4 hours per annotator) of working time. After QA pairs were written, we conducted a verification step to make sure QA pairs were generated correctly. Each QA pair was verified by two other annotators. In each HIT, one predicate (highlighted in one sentence) plus the QA pairs for this predicate were shown (Figure 7). Verification typically takes less than 1 minute to finish, i.e., 3800 * 2 annotators * 1 min = 7600 minutes = 127 hours.
B.2 QA Presence Annotation
Same as QA generation, we collected QA presence labels on MTurk with 27 trained annotators. Presence labels were collected for 50 examples * 10 systems. And each judgment was provided independently by 3 annotators. In each HIT, we showed one system summary and the QA pairs for one predicate in the reference summary (Figure 8). It usually takes less than 2 minutes to finish one HIT. So, in total, 50 examples * 10 systems * 3 annotators * 7.6 predicates * 2 mins = 22800 minutes = 380 hours.
Appendix C Examples, Prompts, and Annotation UIs
@@
Reference: Loeb says he filed the lawsuit and doesn’t want want money from his ”ex” Nick Loeb reportedly wants to prevent Vergara from destroying the embryos . Vergara spoke of freezing embryos with Loeb in a 2013 interview .
Summary (generated by MatchSum): The 42-year-old actress and star of the hit TV sitcom ”Modern Family” split from businessman Nick Loeb in May 2014. Loeb is suing the Colombian-born actress in Los Angeles to prevent Vergara from destroying their two embryos conceived through in vitro fertilization in November 2013, according to published reports by New York Daily News and In Touch magazine.
@@
What did someone say? he filed the lawsuit and does n’t want want money from his ” ex ”
@@
Who doesn’t someone want something from? his ” ex ”
@@
QA to Statement Convert the question and answer into a statement. Start your answer with ”Statement:”